# Assessing the Reasoning Capabilities of LLMs in the context of Evidence-based Claim Verification

### v3

@@ 9 Conclusion
Dataset Selection and Reasoning Types
@@ Appendix A Logical Reasoning Examples
Deductive reasoning or top-down logic is a logical reasoning process where we use inference rules such as modus ponens to deduce the veracity of a conclusion based on multiple hypotheses.
A core element of deductive inference is that if the premises are true, then the conclusion is true. In formal logic, the rules of deduction are infinite Morishita et al. (2023), where the most common ones are modus ponens, syllogism, and elimination. The reader is referred to the works of Morishita et al. (2023) and Saparov et al. (2023) for a more in-depth discussion of deduction rules.
@@ Appendix C Sampling details
[…]
@@ Appendix F Prompts
Appendix G Qualitative Analysis

### v4

@@ 10 Conclusion
Dataset Selection and Reasoning Types.
@@
Acknowledgments
This work was supported by a UKRI/EPSRC Turing AI Fellowship to Maria Liakata (grant ref EP/V030302/1) and the Alan Turing Institute (grant ref EP/N510129/1). This work was also supported by the Engineering and Physical Sciences Research Council [grant number EP/Y009800/1], through funding from Responsible Ai UK (KP0016) as a Keystone project lead by Maria Liakata. F. Ruggeri is partially supported by the project European Commission’s NextGeneration EU programme, PNRR – M4C2 – Investimento 1.3, Partenariato Esteso, PE00000013 - “FAIR - Future Artificial Intelligence Research” – Spoke 8 “Pervasive AI” and by the European Union’s Justice Programme under Grant Agreement No. 101087342 for the project “Principles Of Law In National and European VAT”. Yulan He was supported by the UK Engineering and Physical Sciences Research Council through a Turing AI Fellowship (grant no. EP/V020579/1, EP/V020579/2).
@@ Appendix A Logical Reasoning Examples
Deductive reasoning or top-down logic is a logical reasoning process where we use inference rules such as modus ponens to deduce the veracity of a conclusion based on multiple hypotheses. A core element of deductive inference is that if the premises are true, then the conclusion is true. In formal logic, the rules of deduction are infinite Morishita et al. (2023), where the most common ones are modus ponens, syllogism, and elimination. The reader is referred to the works of Morishita et al. (2023) and Saparov et al. (2023) for a more in-depth discussion of deduction rules.
@@ Appendix C Sampling details
Figure A1 shows the distribution of cosine similarity (denoted as Sim Score), BERTScore and BLEURT score. We used this distribution to set up two different thresholds for deductive and abductive samples. We found that the Bertscores and Sim Scores for abductive sample did not seem to overlap. However, the deductive score had overlap between them. From this observation, we derived the following thresholds.
@@ Appendix F Prompts
Appendix G Human Rationale Evaluation
A third annotator judges the quality of the rationales based on coherence, relevance, and consistency. We denote this annotator as evaluator for readability. The evaluator only considers rationales where annotators agree with each other. In particular, the evaluator selects rationales on the basis of containing the most relevant and coherent information, and providing a consistent narrative. For samples where annotators disagree, the rationale of the annotator with most rationales within the agreement set was considered.
Appendix H Qualitative Analysis
@@
Appendix I Statistical Significance Tests
We run a non-parametric Mann Whitney U test McKnight and Najab (2010) as it has no normality assumption and works with unequal population sizes. Table A12 reports results. Our hypotheses are as follows.
Null Hypothesis ($H_{0}$).
There’s no difference between the distributions in accuracy when predicting veracity for deductive and abductive reasoning.
Alternative Hypothesis ($H_{1}$).
There is a difference — the model behaves differently on abductive reasoning cases.
Model VitaminC CLIMATE-FEVER PHEMEPlus
Claude ZS $7.19e^{-4}$ $1.35e^{-12}$ $2.61e^{-3}$
Claude ZS CoT $2.42e^{-3}$ $7.32e^{-11}$ $5.35e^{-3}$
Claude M-CoT $2.45e^{-4}$ $1.00e^{-17}$ $1.51e^{-2}$
GPT-4 ZS $4.44e^{-5}$ $1.78e^{-11}$ $1.31e^{-5}$
GPT-4 ZS CoT $2.08e^{-5}$ $4.20e^{-12}$ $1.24e^{-4}$
GPT-4 M-CoT $3.25e^{-6}$ $1.39e^{-8}$ $4.14e^{-5}$
GPT-4o ZS $3.75e^{-4}$ $7.56e^{-10}$ $2.45e^{-4}$
GPT-4o ZS CoT $4.63e^{-4}$ $7.32e^{-11}$ $3.75e^{-4}$
GPT-4o M-CoT $4.35e^{-6}$ $5.08e^{-10}$ $8.65e^{-5}$
Table A12: Two-sided Mann Whitney U test results.